Instructions for Coling 2008 Proceedings

First Author Affiliation / Address line 1 Affiliation / Address line 2 Affiliation / Address line 3

email@domain

Second Author Affiliation / Address line 1 Affiliation / Address line 2 Affiliation / Address line 3

        email@domain

Abstract

This document contains the instructions for preparing a camera-ready manuscript for the proceedings of Coling 2008. The document itself conforms to its own specifications, and is therefore an example of what your manuscript should look like. Authors are asked to conform to all the directions reported in this document.

1 Credits

This document has been adapted from the instructions for ACL-07, adapted from those for ACL/Coling-06, which were in turn adapted from those for EACL-06, which were in turn adapted from the instructions for ACL-05 and EACL-03. All these were based on the formats of earlier ACL and EACL Conference proceedings. Those versions were written by several people, including John Chen, Henry S. Thompson and Donald Walker.

2 Introduction

The following formatting instructions are directed to authors of papers accepted for publication in Coling 2008 proceedings, including the main conference, workshops, posters and demos. The conference website http://www.coling2008.org.uk has additional submission information for them. All authors are required to adhere to these specifications. Authors are required to provide a Portable Document Format (PDF) of their papers. The proceedings will be printed on A4 paper. Authors from countries in which access to word processing systems is limited should contact the publications chair Roger Evans (R.P.Evans@brighton.ac.uk) as soon as possible.

3 General Instructions

Manuscripts must be in two-column format. Exceptions to the two-column format include the title, authors’ names and complete addresses, which must be centered at the top of the first page, and any full-width figures or tables (see the guidelines in Subsection 3.5). Type single-spaced. Start all pages directly under the top margin. See the guidelines later regarding formatting the first page.

Unless otherwise specified, the maximum length of a manuscript is 8 pages for regular and workshop papers and 4 pages for posters and demos, printed single-sided (see Section 4 for additional information on the maximum number of pages).

3.1 Electronically-available resources

This description is provided in L^{A}TEX2e (coling08.tex) along with the L^{A}TEX2e style file used to format it (coling08.sty) and a Coling bibliography style (coling.bst); and in PDF format (coling08.pdf). These files are all available at http://www.coling08.org.uk/style.html. There is also a Microsoft Word document template (coling08.dot) available at the same URL. We strongly recommend the use of these style files, which have been appropriately tailored for the Coling 2008 proceedings.

3.2 Format of Electronic Manuscript

For the production of the electronic manuscript you must use Adobe’s Portable Document Format (PDF). This format can be generated from postscript files. On Unix systems, you can use ps2pdf for this purpose. Under Microsoft Windows, you can use Adobe’s Distiller or GSview (File>Convert>pdfwrite); if you have cygwin installed, you can use dvipdf or ps2pdf. Note that some word processing programs generate PDF which may not include all the necessary fonts (esp. tree diagrams, symbols). When you print or create the PDF file, there is usually an option in your printer setup to include none, all or just non-standard fonts. Please make sure that you select the option of including ALL the fonts. Before sending it, test your PDF by printing it from a computer different from the one where it was created. Moreover, some word processors may generate very large postscript / PDF files, where each page is rendered as an image. Such images may reproduce poorly. In this case, try alternative ways to obtain the postscript and / or PDF. One way on some systems is to install a driver for a postscript printer, send your document to the printer specifying “Output to a file”, then convert the file to PDF.

It is of utmost importance to specify the A4 format (21.0 cm x 29.7 cm) / (8.3 in x 11.7 in) when formatting the paper. When working with dvips, for instance, one should specify -t a4.

Print-outs of the PDF file on A4 paper should look like the present document, which conforms to the formatting requirements. Note that in order for your paper to print correctly, you should disable centering and scale-to-fit options on your printer. If you cannot meet the above requirements about the production of your camera-ready paper, please contact the publications chair as soon as possible.

3.3 Layout

Format manuscripts two columns to a page, in the manner these instructions are formatted. The exact dimensions for a page on A4 paper are:

• Left and right margins: 2.5 cm (1 in)

• Top margin: 2.5 cm (1 in)

• Bottom margin: 2.5 cm (1 in)

• Column width: 7.7 cm (3.1 in)

• Column height: 23.7 cm (9.4 in)

• Gap between columns: 0.5 cm (0.2 in)

3.4 Fonts

For uniformity, Adobe’s Times Roman font should be used. In L^{A}TEX2e this is accomplished by putting

    \usepackage{times}
    \usepackage{latexsym}

in the preamble. If Times Roman is unavailable, use Computer Modern Roman (L^{A}TEX2e’s default). Note that the latter is about 10% less dense than Adobe’s Times Roman font.

3.5 The First Page

Center the title, authors’ names and affiliations across both columns. Do not use footnotes for affiliations. Do not include the paper ID number assigned during the submission process. Use the two-column format only when you begin the abstract.

Title: Place the title centered at the top of the first page, in a 15-point bold font. Long titles should be typed on two lines without a blank line intervening. Approximately, put the title at 2.5 cm (1 in) from the top of the page, followed by a blank line, then the authors’ names, and the affiliations on the following line. Do not use only initials for given names (middle initials are allowed). Avoid capitalizing last names. The affiliation should contain the author’s complete address, and if possible an electronic mail address. Leave about 2 cm (0.75 in) between the affiliation and the body of the first page.

Abstract: Type the abstract at the beginning of the first column. The width of the abstract text should be smaller than the width of the columns for the text in the body of the paper by about 0.6 cm (0.25 in) on each side. Center the word Abstract in a 12 point bold font above the body of the abstract. The abstract should be a concise summary of the general thesis and conclusions of the paper. It should be no longer than 200 words.

Text: Begin typing the main body of the text immediately after the abstract, observing the twocolumn format as shown in the present document. Use 11 point font for text. Indent when starting a new paragraph, except for the first paragraph following each heading.

3.6 Sections

Headings: Type and label section and subsection headings in the style shown on the present document. Use numbered sections (Arabic numerals) in order to facilitate cross references. Number subsections with the section number and the subsection number separated by a dot, in Arabic numerals. Do not number subsubsections. Use 11 point font for subsection headings and 12 point font for section headings.

Citations: Citations within the text appear in parentheses as (Gusfield, 1997) or, if the author’s name appears in the text itself, as Gusfield (1997). Citations in parentheses should not be used as linguistic phrases; for example, instead of “(Gusfield, 1997) argues that … ” use “Gusfield (1997) argues that … ”. Append lowercase letters to the year in cases of ambiguity. Treat double authors as in (Aho and Ullman, 1972), but write as in (Chandra et al., 1981) when more than two authors are involved. Collapse multiple citations in parenthesis as in (Gusfield, 1997; Aho and Ullman, 1972).

References: Gather the full set of references together under the heading References; place the section before any Appendices, unless they contain references. Arrange the references alphabetically by the first author’s last-name, rather than by order of occurrence in the text, and invert the first-name and last-name of the first author (only). Provide as complete a citation as possible, using a consistent format, such as the one for Computational Linguistics or the one in the Publication Manual of the American Psychological Association (American Psychological Association, 1983). Use of full names for authors rather than initials is preferred. Use full names for journals and conferences, not abbreviations (for example “45th Meeting of the Association for Computational Linguistics”, not “ACL07”).

The L^{A}TEX2e and BibTEX style files provided roughly fit the American Psychological Association format, allowing regular citations, short citations and multiple citations as described above.

Appendices: Appendices, if any, directly follow the text and the references (but see above). Letter them in sequence and provide an informative title: Appendix A. Title of Appendix.

Acknowledgement section should go as a last section immediately before the references. Do not number the acknowledgement section.

3.7 Footnotes

Footnotes: Put footnotes at the bottom of the page and use 9-point font. They may be numbered or referred to by asterisks or other symbols.^{1} Footnotes should be separated from the main text by a line.^{2}

3.8 Graphics

Illustrations: Place figures, tables, and photographs in the paper near where they are first discussed, rather than at the end, if possible. Wide illustrations may run across both columns. Do not use color illustrations as they may reproduce poorly.

Captions: Provide a caption for every illustration; number each one sequentially in the form: “Figure 1. Caption of the Figure.”, “Table 1. Caption of the Table.” Type the captions of the figures and tables below the body, using 11-point text.

4 Length of Submission

Unless otherwise specified, the maximum length is 8 pages for regular and workshop papers and 4 pages for posters and demos. The page limit should be observed strictly. All illustrations, references, and appendices must be accommodated within these page limits, following the formatting instructions given in the present document.

References

Aho, Alfred V. and Jeffrey D. Ullman. 1972. The Theory of Parsing, Translation and Compiling, volume 1. Prentice-Hall, Englewood Cliffs, NJ.

American Psychological Association. 1983. Publications Manual. American Psychological Association, Washington, DC.

Association for Computing Machinery. 1983. Computing Reviews, 24(11):503–512.

Chandra, Ashok K., Dexter C. Kozen, and Larry J. Stockmeyer. 1981. Alternation. Journal of the Association for Computing Machinery, 28(1):114–133.

Gusfield, Dan. 1997. Algorithms on Strings, Trees and Sequences. Cambridge University Press, Cambridge, UK.

1This is how a footnote should appear.

2Note the line separating the footnotes from the text.